CC12:

I need math help please A. 24 B. 16 C. 8 D. 4

CC12:

1 attachment
CC12:

I think the answer is 24

jhonyy9:

AC = AE + EC

jhonyy9:

AC = x^2 -8 +2x x^2 +2x -8 = 0 (x+4)(x-2)=0

CC12:

8

jhonyy9:

first x+4=0 x=-4 and x-2=0 x=2

CC12:

so its 8 or 24

CC12:

ok nvm its 4 right

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 AC = x^2 -8 +2x x^2 -8 = 2x x^2 -2x -8 = 0 (x-4)(x+2)=0 \(\color{#0cbb34}{\text{End of Quote}}\)

CC12:

Ok thank you so much

jhonyy9:

np my pleasure